OpenStudy (anonymous):

Solve "6 cos (2x) - 2 sin (x) = 4" for [0,2Pi)

OpenStudy (dumbcow):

double angle identity: cos(2x) = 1-2sin^2

OpenStudy (dumbcow):

treat it like a quadratic equation, set equal to 0 --> 12sin^2 +2sin -2 = 0

OpenStudy (anonymous):

I never thought of that! Hold on

OpenStudy (anonymous):

sin x = y 12y^2 + 2y - 2 = 0 (4y + 2)(3y - 1) = 0 y = -1/2 or y = 1/3

OpenStudy (dumbcow):

very good, not use inverse sine to get x values

OpenStudy (dumbcow):

now*

OpenStudy (anonymous):

sin x = -1/2 x = -.523598 sin x = 1/3 x = .339836

OpenStudy (anonymous):

Would I have to turn this into a "better looking" value? Such as pi over a number?

OpenStudy (dumbcow):

yes but remember x is in [0,2pi] make it positive angle also use unit circle, there should be 4 total solutions

OpenStudy (anonymous):

Right, how would you go about doing that?

OpenStudy (dumbcow):

for sin x = -1/2, yes its x = -pi/3 but for the other one there is no "nice" value

OpenStudy (dumbcow):

 find the angles